# DELIFT: Data Efficient Language model Instruction Fine-Tuning

### v2

@@ Abstract
Fine-tuning large language models (LLMs) is essential for enhancing their performance on specific tasks but is often resource-intensive due to redundant or uninformative data. To address this inefficiency, we introduce DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel algorithm that systematically optimizes data selection across the three key stages of fine-tuning: (1) instruction tuning, (2) task-specific fine-tuning (e.g., reasoning, question-answering), and (3) continual fine-tuning (e.g., incorporating new data versions). Unlike existing methods that focus on single-stage optimization or rely on computationally intensive gradient calculations, DELIFT operates efficiently across all stages. Central to our approach is a pairwise utility metric that quantifies how beneficial a data sample is for improving the model’s responses to other samples, effectively measuring the informational value relative to the model’s current capabilities. By leveraging different submodular functions applied to this metric, DELIFT selects diverse and optimal subsets that are useful across all stages of fine-tuning. Experiments across various tasks and model scales demonstrate that DELIFT can reduce the fine-tuning data size by up to 70% without compromising performance, offering significant computational savings and outperforming existing methods in both efficiency and efficacy.
@@ 1 Introduction
Fine-tuning large language models (LLMs) is pivotal for adapting these powerful architectures (Devlin et al., 2019; Brown et al., 2020a; Touvron et al., 2023) to specialized tasks such as intricate reasoning, precise question-answering, and the seamless integration of new information (Ouyang et al., 2022). This transformation—from a general-purpose model to a task-specific agent—heavily relies on the quality and nature of the data employed during fine-tuning, which critically determines the model’s subsequent performance (Wei et al., 2022; Zhou et al., 2023; Hoffmann et al., 2024).
The effectiveness of fine-tuning hinges on the quality, diversity, and relevance of the selected data (Gururangan et al., 2020; Wei et al., 2022; Zhou et al., 2023). High-quality data ensures accurate learning, diverse data enhances generalization, and relevant data aligns the model’s capabilities with specific application needs. However, optimizing data selection across different fine-tuning phases remains a significant challenge, leading to our central research question:
@@
How can we create a unified framework for efficient data selection across all fine-tuning stages of LLMs, while optimizing performance and maximizing data efficiency?
To address this challenge, we present DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel, unified, and computationally efficient algorithm engineered to optimize data selection across all stages of the fine-tuning process. The key innovation of DELIFT lies in its pairwise utility metric, which assesses the informational value of data samples relative to both the model’s current capabilities and other samples within the dataset. This metric, combined with submodular optimization techniques, allows DELIFT to efficiently select optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs.
The typical fine-tuning process comprises three key stages: 1. Instruction Tuning: Enhances the model’s ability to follow general instructions (Mishra et al., 2022; Wei et al., 2022; Longpre et al., 2023); 2. Task-Specific Fine-Tuning: Refines the model’s expertise in specific domains (Gururangan et al., 2020; Cobbe et al., 2021); 3. Continual Fine-tuning: Enables the model to integrate new information while mitigating catastrophic forgetting (Madotto et al., 2021; Wu et al., 2024). DELIFT is able to optimize data selection processes across all three stages. Additionally, DELIFT offers significant benefits for In-Context Learning (ICL) (Brown et al., 2020b; Xue et al., 2024). By utilizing the selected subsets as the ICL example pool, DELIFT achieves similar or better performance compared to using the entire dataset, thereby enhancing data efficiency in ICL scenarios. This dual functionality is empirically validated in our experimental results.
Existing data selection methodologies often fail to address the nuanced requirements of the aforementioned distinct fine-tuning stages. Many approaches are tailored to a single stage, lacking the adaptability needed for comprehensive fine-tuning (Xia et al., 2024; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024). Others depend on computationally intensive procedures, such as exhaustive gradient computations, rendering them impractical for large-scale models and datasets (Killamsetty et al., 2021b; a; Xia et al., 2024; Zhang et al., 2024). Additionally, some methods utilize features obtained from an independent model that are not specifically aligned with the model undergoing fine-tuning, reducing their effectiveness (Killamsetty et al., 2023; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024; Du et al., 2023).
DELIFT addresses these limitations by adapting to the unique requirements of each fine-tuning stage. 1. Instruction Tuning: Selects diverse data to enhance general instruction-following capabilities; 2. Task-Specific Fine-Tuning: Prioritizes data that is aligned with the target task, to refine specialized expertise; 3. Continual Fine-tuning: Identifies novel, complementary information to expand the model’s knowledge base while safeguarding against catastrophic forgetting.
Figure 1 illustrates how DELIFT optimizes data selection across these stages, demonstrating the selection and pruning processes in each fine-tuning phase. By leveraging submodular optimization techniques (Fujishige, 2005; Bilmes, 2022) and submodular information measures (Iyer et al., 2021), DELIFT efficiently selects optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs. This approach effectively balances data utility and computational efficiency.
@@
Our key contributions are as follows:
1) Versatile Pairwise Utility Metric: A novel, easy-to-compute metric for assessing data informativeness, incorporating model feedback applicable across all fine-tuning stages.
2) Unified Data Selection Algorithm: DELIFT systematically optimizes data selection for instruction tuning, task-specific fine-tuning, and continual fine-tuning within a single framework.
3) Computational Efficiency: Circumvents resource-intensive operations, ensuring scalability to large datasets and models. DELIFT achieves at least 70% reduction in computational time compared to gradient-based methods on benchmark tasks.
4) Enhanced Performance with Reduced Data: Demonstrates the ability to reduce fine-tuning data size by up to 70% without compromising performance, and achieves comparable efficacy as to utilizing the full dataset.
5) Improvement over Existing Methods: Outperforms current data selection techniques by up to 26% in effectiveness across diverse tasks and model scales (see Section  4).
The remainder of this paper is organized as follows: Section 2 provides background on fine-tuning LLMs and reviews related work. Section 3 details the methodology behind DELIFT, including the development of our pairwise utility metric and the submodular optimization process. Section 4 presents experimental results that showcase the effectiveness and efficiency of our method. Section 5 discusses the implications of our findings and potential future directions. Finally, we release our code base for further research.
@@ 2 Related Work
Efficient data subset selection is vital for enhancing training efficiency in deep neural networks while maintaining or improving model performance. This section categorizes existing subset selection methods into model-independent and model-dependent approaches and identifies the gaps our work addresses. Model-independent subset selection methods focus on selecting representative subsets without model-specific feedback. Common approaches include using pre-trained sentence embeddings with distance or clustering metrics (Bukharin & Zhao, 2024; Sorscher et al., 2023; Killamsetty et al., 2023; Du et al., 2023; Bhatt et al., 2024), as well as employing large models like GPT-4 or pre-trained reward models for high-quality data filtering (Du et al., 2023; Chen et al., 2024). However, these methods often struggle to translate the assessed diversity or quality into downstream utility. Model-dependent subset selection aims to identify data samples beneficial to the downstream model by analyzing features like per-sample gradients or loss values. Methods such as GradMatch (Killamsetty et al., 2021a), CRAIG (Mirzasoleiman et al., 2020), and TAGCOS (Zhang et al., 2024) focus on selecting samples that approximate the gradient updates over the full dataset. GLISTER (Killamsetty et al., 2021b) employs bilevel optimization to align gradients from selected subsets with those of a validation set. LESS (Xia et al., 2024) proposes computing gradients through LoRA fine-tuning to reduce the computational cost of gradient computation and utilizes random projection to address gradient dimensionality issues. Li et al. (2024) proposed the IFD score, a computationally efficient model-dependent metric that assesses instruction difficulty to filter challenging samples, though it does not guarantee data diversity. While effective in capturing useful samples, these methods often face computational challenges, especially with LLMs. Persistent limitations across these methods include: (i) Limited Adaptability across different fine-tuning stages, (ii) Computational Intensity due to model feedback reliance, (iii) Lack of Unified Solutions applicable across all fine-tuning phases, and (iv) ineffective Redundancy Handling. DELIFT addresses these limitations through a novel pairwise utility metric, which effectively aligns with the model’s evolving capabilities throughout fine-tuning. By integrating submodular optimization with pairwise model-dependent metrics that evaluate relative sample utility, DELIFT minimizes redundancy while maximizing adaptability and computational efficiency. This approach proves effective across diverse use cases including instruction tuning, task-specific fine-tuning, continual fine-tuning, and In-Context Learning (ICL), offering a versatile and scalable solution for data subset selection.
@@ 3 Methodology
This section presents foundational concepts and the specific approach of DELIFT, focusing on data subset selection through a utility-based kernel integrated with submodular optimization techniques.
3.1 Notation
Let $\mathcal{D}$ denote the fine-tuning dataset, comprising elements $d_{i}=(x_{i},y_{i})$, where $x_{i}$ is the input sequence and $y_{i}$ is the corresponding output sequence. Our objective is to select a subset $\mathcal{A}\subseteq\mathcal{D}$ that maximizes the model’s performance while minimizing computational costs. The selection strategy adapts based on the fine-tuning objective, which may include instruction tuning, task-specific adaptation, or continual learning.
3.2 Utility-Based Kernel
At the core of DELIFT lies the utility-based kernel, a mechanism designed to quantify the informativeness of one data point when used as an in-context example for another. Consider two data points, $(x_{i},y_{i})$ and $(x_{j},y_{j})$. The utility of data point $j$ relative to data point $i$, denoted as $UF_{ij}$, is defined as:
$UF_{ij}=d(GT_{i},p(y_{i}\mid x_{i}))-d(GT_{i},p(y_{i}\mid x_{j},y_{j},x_{i})),$ (1)
where: $d(\cdot,\cdot)$ is a length-normalized distance metric between two probability distributions, $GT_{i}$ is the ground truth distribution for the sequence $y_{i}$, modeled as a vector of ones for each token to signify perfect prediction, $p(y_{i}\mid x_{i})$ is the model’s predicted probability distribution for $y_{i}$ given only the input $x_{i}$, $p(y_{i}\mid x_{j},y_{j},x_{i})$ is the predicted distribution for $y_{i}$ when the model is provided with $(x_{j},y_{j})$ as an in-context example followed by $x_{i}$.
The distance metric $d(p,q)$ is calculated using the length normalized L2 norm and is defined as:
$d(p,q)=\sqrt{\frac{\sum_{k=1}^{N}(p_{k}-q_{k})^{2}}{N}},$ (2)
where $p_{k}$ and $q_{k}$ are the $k$-th elements of the flattened probability distributions $p$ and $q$, respectively. Importantly, $N$ varies with each data sample and corresponds to the number of tokens in the ground truth response $y_{i}$. This normalization ensures that the distance measure remains scale-invariant across different sequence lengths and vocabulary sizes.
To compute the probability distributions accurately, we employ the teacher forcing technique (Williams & Zipser, 1989). This method ensures that the model uses the ground truth of previous tokens when predicting each subsequent token in the sequence, enabling reliable measurement of prediction accuracy.
The utility value $UF_{ij}$ measures the improvement in prediction accuracy for the entire sequence $(x_{i},y_{i})$ when utilizing $(x_{j},y_{j})$ as an in-context example. A positive $UF_{ij}$ indicates that including data point $j$ enhances the model’s prediction accuracy for $i$, whereas a negative value suggests an adverse effect.
3.3 Submodular Functions for Dataset Selection
To optimize the selection of informative data subsets, DELIFT leverages submodular functions (Fujishige, 2005). Submodular functions are characterized by the property of diminishing marginal returns, making them ideal for selecting diverse, informative, and non-redundant subsets. Submodular function maximization can be efficiently approximated using a greedy algorithm, with a provable approximation guarantee of $1-\frac{1}{e}$ of the optimal solution (Nemhauser et al., 1978).
We employ three tailored submodular functions (Iyer et al., 2021), each suited to a specific fine-tuning stage:
3.3.1 Facility Location (FL)
From an information perspective, the Facility Location function maximizes the coverage of the information space. It ensures that the selected subset $\mathcal{A}$ contains examples that are collectively representative of the entire dataset’s $\mathcal{D}$ information content. This is particularly useful in instruction tuning, where we aim to capture a diverse range of instruction types and their informational content. It is defined as the following where $s_{ij}$ is the similarity measure between data points $i$ and $j$:
$f_{FL}(\mathcal{A})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij},$ (3)
3.3.2 Facility Location Mutual Information (FLMI)
The FLMI function is designed to maximize the mutual information between the selected subset $\mathcal{A}$ and the target domain dataset $\mathcal{D_{T}}$. In our context, it ensures that the selected data points are not only informative in general but also particularly relevant to the specific task at hand. This makes it ideal for task-specific fine-tuning, where we want to bridge the gap between general knowledge and task-specific information. It is defined below where $\eta$ is a scaling factor (set to $1$ in our experiments):
$f_{FLMI}(\mathcal{A};\mathcal{D}_{T})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}+\eta\sum_{j\in\mathcal{A}}\max_{i\in\mathcal{D}_{T}}s_{ij},$ (4)
@@ 3.3.2 Facility Location Mutual Information (FLMI)
3.3.3 Facility Location Conditional Gain (FLCG)
From an information-theoretic standpoint, the FLCG function aims to maximize the conditional information gain of the selected subset $\mathcal{A}$ given the existing dataset $\mathcal{D_{E}}$. It quantifies how much new information each data point brings, conditional on what the model already knows. This is crucial for continual fine-tuning, where we want to avoid redundancy and focus on novel, complementary information that expands the model’s knowledge base without unnecessary repetition. It is defined as the following where $\nu$ is a scaling factor (set to $1$ in our experiments).
$f_{FLCG}(\mathcal{A}\mid\mathcal{D}_{E})=\sum_{i\in\mathcal{D}}\max\left(\max_{j\in\mathcal{A}}s_{ij}-\nu\max_{k\in\mathcal{D}_{E}}s_{ik},0\right),$ (5)
Each submodular function, when combined with our utility-based kernel, guides the selection of data subsets tailored to the specific fine-tuning stage. This ensures that DELIFT selects the most informative and diverse examples, maximizing the efficiency and effectiveness of fine-tuning.
3.4 Utility Kernel as Feature Space
Our approach utilizes the utility-based kernel as a feature space for data selection, representing a significant departure from traditional semantic similarity-based methods. Traditional methods often rely on sentence embeddings (SE) to capture static semantic similarities between data points. In contrast, our utility-based kernel measures the actual impact of examples on model performance, providing a dynamic and task-specific assessment.
This distinction is crucial for two main reasons: 1. Semantic Diversity vs. Performance Enhancement: While SE-based methods select diverse examples solely based on semantic content, our utility-based approach selects examples that demonstrably improve model performance across various inputs; 2. Model-Aware Selection: The utility-based kernel is attuned to the model’s current capabilities and weaknesses, enabling the selection of data points that are most beneficial for enhancing performance on the target task.
By integrating the utility-based kernel with the aforementioned submodular functions DELIFT tailors the data selection process to each fine-tuning stage: instruction tuning, task-specific fine-tuning, and continual learning.
3.5 Data Subset Selection Algorithm
To operationalize our data selection approach, we employ a **greedy algorithm** that iteratively builds the subset $\mathcal{A}$ by selecting the data point that offers the maximum marginal gain in the chosen submodular function.
@@ 3.5 Data Subset Selection Algorithm
This greedy algorithm ensures that each addition to the subset $\mathcal{A}$ maximizes the marginal gain in the submodular function $f$. By iteratively selecting the most beneficial data points according to the utility-based kernel and the specific submodular function tailored to the fine-tuning stage, DELIFT efficiently utilizes the data budget to select the most informative examples.
The complete subset selection process involves the following steps: 1. Compute the Utility-Based Kernel: Calculate $UF_{ij}$ for all relevant pairs of data points in the dataset to assess their informativeness; 2. Select the Appropriate Submodular Function: Depending on the fine-tuning stage (instruction tuning, task-specific fine-tuning, or continual fine-tuning), choose the corresponding submodular function (FL, FLMI, or FLCG); 3. Apply the Greedy Maximization Algorithm: Use Algorithm 1 to iteratively build the subset $\mathcal{A}$ by selecting data points that offer the highest marginal gain according to the selected submodular function.
By synergizing our novel utility-based kernel with submodular optimization, DELIFT achieves data-efficient fine-tuning that effectively addresses both redundancy and informativeness in the data selection process, optimizing the model’s performance across various tasks and domains.

### v3

@@ Abstract
Fine-tuning large language models (LLMs) is crucial for task specialization but often becomes resource-intensive due to redundant or uninformative data. Existing data selection methods typically rely either on computationally expensive gradient-based metrics or static embeddings that fail to adapt dynamically to the model’s evolving state, thus limiting their practical effectiveness. To address this, we propose DELIFT (Data Efficient Language model Instruction Fine-Tuning), leveraging a novel, computationally efficient utility metric inspired by In-Context Learning (ICL). Our ICL-based metric measures the informational value of each data sample by quantifying its effectiveness as an in-context example in improving model predictions for other samples, reflecting its actual contribution relative to the model’s current state. Integrated with tailored submodular optimization methods, DELIFT systematically selects diverse, informative subsets optimized specifically for each fine-tuning stage: instruction tuning, task-specific adaptation, and continual fine-tuning. Experimental results across multiple datasets and model scales show DELIFT reduces fine-tuning data requirements by up to 70% without compromising performance, consistently outperforming existing methods by up to 26% in effectiveness and efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have become indispensable for solving a variety of natural language processing tasks, ranging from question answering and summarization to complex dialogue and reasoning (Brown et al., 2020; Touvron et al., 2023). Despite their remarkable adaptability, fine-tuning LLMs often requires enormous computational resources and time, especially when a significant portion of the training data is either redundant or uninformative (Gururangan et al., 2020; Sorscher et al., 2023). This challenge grows more critical with increasing model and dataset sizes, posing a key limitation to the broader deployment of LLMs.
Existing data selection methods generally fall under two paradigms: (1) static embedding-based approaches that compute sample similarities without reflecting the model’s evolving state (Bukharin & Zhao, 2024; Chen et al., 2024), and (2) gradient-based methods that offer more model-specific feedback but often entail prohibitive computational overhead, especially for large-scale models (Killamsetty et al., 2021b; Xia et al., 2024). Although both paradigms can yield initial benefits, they often fail to account for how a model’s knowledge shifts over multiple fine-tuning phases: (1) Instruction Tuning (Mishra et al., 2022; Wei et al., 2022; Longpre et al., 2023), which enhances the model’s ability to follow diverse instructions; (2) Task-Specific Fine-Tuning (Gururangan et al., 2020; Cobbe et al., 2021), which focuses on refining domain expertise; and (3) Continual Fine-Tuning (Madotto et al., 2021; Wu et al., 2024), which incrementally incorporates new knowledge while mitigating catastrophic forgetting.
Thus, a natural question arises:
@@
Can we develop a unified, computationally efficient data selection framework that adapts to all stages of fine-tuning and maximizes model performance while minimizing data redundancy?
@@
In this paper, we introduce DELIFT (Data-Efficient Language Model Instruction Fine-Tuning), a single-stop solution designed to address data selection across all fine-tuning stages within a single framework. DELIFT is grounded in information theory yet uses the practical intuition of in-context examples to assess the ’information gain’ of each data sample relative to the current state of a model. Specifically, we propose a new utility metric that captures how effectively one sample improves the model’s prediction of another. By combining these pairwise utilities with submodular optimization, DELIFT generates diverse, nonredundant subsets uniquely tailored to each fine-tuning phase.
We evaluated DELIFT on various tasks and model scales, consistently observing that it can prune up to 70% of the training data without hurting performance - and often improving it - outperforming existing methods by up to 26% in efficiency and effectiveness. In doing so, we show that careful, utility-driven data selection can be far more effective than sheer data volume, opening the door to more resource-friendly and targeted fine-tuning.
Our primary contributions are as follows.
1. A unified information-theoretic data selection paradigm that leverages pairwise utilities grounded in conditional pointwise mutual information, making it adaptable to instruction tuning, task-specific adaptation, and continual fine-tuning.
2. A single-stop, submodular optimization framework that integrates these utilities to provide diverse, high-value subsets for each fine-tuning stage without incurring prohibitive computation.
3. Extensive empirical validation showing up to 70% data reduction with minimal (and sometimes zero) performance loss across multiple domains, demonstrating substantial gains in both efficacy and efficiency.
The remainder of this paper is organized as follows. Section 2 reviews prior work on data-efficient strategies for fine-tuning LLMs and situates our approach within the literature. Section 3 introduces our information-theoretic utility metric and describes how it integrates with submodular optimization to enable data selection across diverse fine-tuning stages. Section 4 presents comprehensive experiments demonstrating the effectiveness and efficiency of our framework on multiple tasks and models. Finally, Section 5 discusses the broader implications of our results, outlines limitations, and suggests directions for future research. Our complete code base is publicly available at https://github.com/agarwalishika/delift, enabling further exploration and replication.
@@ 2 Related Work
Data Subset Selection for Deep Neural Networks.
Selecting an informative subset of training samples is a longstanding strategy to reduce computational costs and enhance model generalization. Model-Independent Approaches. Traditional model-independent techniques, such as clustering or distance metrics on pre-trained embeddings (Bukharin & Zhao, 2024; Du et al., 2023; Killamsetty et al., 2023), capture broad semantic similarities but do not reflect the model’s changing state, limiting their effectiveness during iterative fine-tuning. Model-Dependent Approaches. Model-dependent methods incorporate the model’s evolving knowledge by analyzing gradients or loss values (Killamsetty et al., 2021b; a; Xia et al., 2024), often outperforming static approaches. However, performing gradient or influence estimations at scale becomes prohibitively expensive for large models. Techniques like LESS (Xia et al., 2024) alleviate some overhead via parameter-efficient fine-tuning (e.g., LoRA), , yet still incur repeated gradient or influence calculations that scale poorly with dataset size. Subset Selection with LLM Feedback. Another emerging direction leverages LLM feedback to score or filter training samples. For instance, SelectIT (Liu et al., 2024) employs self-reflection prompts to rate data quality, while filtering approaches using GPT-4 (Chen et al., 2024) rely on external heuristics. Though these provide a form of model-aware sampling, they typically lack a principled theoretical grounding. In addition, all these approaches primarily target a single fine-tuning stage, limiting their adaptability for instruction tuning, task-specific adaptation, or continual learning.
Our Contribution.
In contrast, we present a unified, information-theoretic framework that operates effectively across all fine-tuning stages: instruction tuning, task-specific adaptation, and continual fine-tuning. Our novel utility metric quantifies how one data point aids the prediction of another, mirroring the model’s evolving knowledge. Integrated within a submodular selection paradigm (Fujishige, 2005; Iyer et al., 2021), this approach balances diversity, coverage, and informativeness throughout the entire fine-tuning pipeline. As a result, we bridge the gap left by existing methods that are either restricted to a single phase or computationally infeasible at scale, demonstrating consistent performance improvements and notable efficiency gains.
@@ 3 Methodology
Our goal is to efficiently identify a subset of data that maximizes the performance of large language models across three fine-tuning stages: (1) Instruction Tuning, (2) Task-Specific Fine-Tuning, and (3) Continual Fine-Tuning. This section first introduces our novel information-theoretic pairwise utility metric (Section 3.1) and then explains how we leverage submodular optimization to achieve data-efficient selection (Section 3.2). Finally, we show how these components unify into a single-stop solution for all fine-tuning stages (Section 3.3).
3.1 Pairwise Utility Metric
Let $\mathcal{D}=\{(x_{i},y_{i})\}$ be a training set, where each $x_{i}$ is an input sequence and $y_{i}$ is the corresponding output. Consider two samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, and let $GT_{i}$ denote the ideal “ground truth” distribution that assigns probability 1 to the token sequence $y_{i}$ and 0 otherwise. We define $p(y_{i}\mid x_{i})$ as the model’s predicted probability distribution of $y_{i}$ given $x_{i}$ alone, and $p(y_{i}\mid x_{i},x_{j},y_{j})$ as the predicted distribution of $y_{i}$ when $(x_{j},y_{j})$ is also provided (e.g., as an in-context example).
Definition (Utility Function).
We capture the information gain of $(x_{j},y_{j})$ for predicting $(x_{i},y_{i})$ via:
$UF_{ij}\;=\;d\bigl{(}GT_{i},\;p(y_{i}\mid x_{i})\bigr{)}\;-\;d\bigl{(}GT_{i},\;p(y_{i}\mid x_{i},x_{j},y_{j})\bigr{)},$ (1)
where $d(\cdot,\cdot)$ is a distance between probability distributions.
Information-Theoretic Interpretation.
Below we state a simplified version of our main theoretical result (see Appendix A for the full proof):
Theorem 1
(Informal Statement) If $d(\cdot,\cdot)$ is chosen to be the Kullback-Leibler (KL) divergence, then the utility $UF_{ij}$ coincides with the (conditional) pointwise mutual information between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$. Formally,
$UF_{ij}\;=\;\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}\;=\;\sum_{t=1}^{T}\log\Bigl{(}\tfrac{p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})}{p(y_{it}\mid x_{i},y_{i,<t})}\Bigr{)}.$
Thus, $UF_{ij}$ captures how much $(x_{j},y_{j})$ informs the prediction of $y_{i}$ given $x_{i}$.
Practical Computation.
In practice, for numerical stability, we adopt a length-normalized Euclidean distance rather than the KL-divergence:
$d\bigl{(}GT_{i},\;p(y_{i}\mid\cdot)\bigr{)}\,=\,\Bigl{\lVert}1-p(y_{i}\mid\cdot)\Bigr{\rVert}_{2},$
where we only extract the probability assigned to each ground-truth token under teacher forcing. This preserves the spirit of the PMI-based formulation while avoiding the instability issues often encountered with near-zero probabilities in large vocabularies. We compute $UF_{ij}$ for all pairs $(i,j)$ once before fine-tuning. Although this step is $O(n^{2})$ in the dataset size, the cost is amortized because the same utility matrix can be reused for different fine-tuning stages or methods.
3.2 Submodular Optimization for Data Selection
After computing $UF_{ij}$, we define a kernel matrix $s_{ij}$ (e.g., set $s_{ij}=\max(UF_{ij},0)$) and utilize it in well-studied submodular functions (Fujishige, 2005; Iyer et al., 2021). Submodularity naturally captures diminishing returns and promotes coverage of diverse yet informative samples.
Objectives.
We primarily adopt three variants:
  1. 1.
    Facility Location (FL): $f_{FL}(\mathcal{A})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}.$
  2. 2.
    Facility Location Mutual Information (FLMI): $f_{FLMI}(\mathcal{A};\mathcal{D}_{T})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}\;+\;\eta\sum_{j\in\mathcal{A}}\max_{i\in\mathcal{D}_{T}}s_{ij}.$
  3. 3.
@@ 3.2 Submodular Optimization for Data Selection
    Facility Location Conditional Gain (FLCG): $f_{FLCG}(\mathcal{A}\mid\mathcal{D}_{E})=\sum_{i\in\mathcal{D}}\max\Bigl{(}\max_{j\in\mathcal{A}}s_{ij}\;-\;\nu\,\max_{k\in\mathcal{D}_{E}}s_{ik},0\Bigr{)}.$
FL supports coverage of $\mathcal{D}$ itself, ideal for instruction tuning; FLMI additionally ties the selection to a task-specific dataset $\mathcal{D}_{T}$; and FLCG incorporates a previously used set $\mathcal{D}_{E}$ to promote continual learning of novel examples.
@@
Subset Selection.
To choose a subset $\mathcal{A}$ of size $k$, we apply a classic greedy heuristic (Nemhauser et al., 1978), each step picking $d^{*}=\arg\max_{d\in\mathcal{D}\setminus\mathcal{A}}[f(\mathcal{A}\cup\{d\})-f(\mathcal{A})]$ (see Algorithm 1). This yields a $1-\tfrac{1}{e}$ approximation factor for submodular functions, ensuring that we obtain near-optimal subsets efficiently.
3.3 A Single-Stop Solution for All Fine-Tuning Stages
By combining the utility-based kernel with the above submodular objectives, we obtain DELIFT, a unified framework that selects data holistically across instruction tuning, task-specific fine-tuning, and continual learning:
  • Instruction Tuning: Use Facility Location (FL) for diverse coverage of general instructions.
  • Task-Specific Fine-Tuning: Use FLMI to align samples with a target benchmark or domain $\mathcal{D}_{T}$.
  • Continual Fine-Tuning: Use FLCG to incorporate new, complementary data while avoiding redundancy with existing knowledge $\mathcal{D}_{E}$.
Despite these distinct objectives, each stage uses the same underlying pairwise utility metric to capture how one sample can enhance the model’s predictions on another. This consistent, model-aware metric is key to DELIFT’s strong performance across diverse tasks and model sizes.